### v1

@@ C.5 Proof of Theorem 1
^$\hat{\nabla}C(K)=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}\hat{C}_{k}U_{k}$
and the true gradient $\nabla C(K)$. {Proof} We bridge this two terms with $\nabla^{\prime}C(K)$ and ¯$\bar{\nabla}C(K))$, where ¯$\bar{\nabla}C(K))$ is defined in Lemma C.3 and $\nabla^{\prime}C(K)$ is defined as:
$\nabla^{\prime}C(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}C^{(l)}(K+U_{k})U_{k}.$
The relation between the estimates and the true gradient is given as:
(77) ^$\hat{\nabla}C(K)-\nabla C(K)=(\hat{\nabla}C(K)-\nabla^{\prime}C(K))+(\nabla^{\prime}C(K)-\bar{\nabla}C(K))+(\bar{\nabla}C(K)-\nabla C(K))$
The third term is analyzed in Lemma C.3. We choose $\epsilon_{nr}:=\epsilon_{n}+\epsilon_{r}$ and $\delta_{n}$, then we can find the exploration radius $r_{\max}$ and $N$ as:
(78a) $r_{\max}$ $=\min\{h(C(K)),\lVert K\rVert,\frac{\epsilon_{r}}{h_{\nabla}(C(K))}\}$
(78b) $N$ $=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{n}^{2}}\left(\alpha_{4}^{2}(C(K),\epsilon_{r})+\frac{\alpha_{5}(C(K),\epsilon_{r})\epsilon_{n}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{n}}\right]$
@@
(79) [¯$\mathbb{P}\left[\lVert\bar{\nabla}C(K)-\nabla C(K)\rVert_{F}\leq\epsilon_{nr}\right]\geq 1-\delta_{n}.$
For the second term, as stated in Lemma C.1, for any desired accuracy $\epsilon_{l}$, we can choose $l$ such that:
(80) $l\geq\frac{2(C(K)+rh_{C}C(K))^{2}}{\epsilon_{l}\lambda_{1}(\Sigma_{w})}\left(\frac{\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{(C(K)+rh_{C}C(K))}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{1}{\lambda_{1}(Q)}\right),$
@@
(81) $|C^{(l)}(\hat{K}_{k})-C(\hat{K}_{k})|\leq\frac{r\epsilon_{l}}{n_{x}n_{u}},\quad\forall k.$
@@
(82) $\lVert\nabla^{\prime}C(K)-\bar{\nabla}C(K)\rVert\leq\left\lVert\frac{1}{n}\sum_{k=1}^{n}\frac{n_{x}n_{u}}{r^{2}}\left[C^{(l)}(\hat{K}_{k})-C(\hat{K}_{k})\right]U_{k}\right\rVert=\epsilon_{l}$
@@
(83) $\lVert x_{t}\rVert\leq\bar{L}=L_{0}+t\bar{w}=L_{0}+t\sqrt{\frac{n_{x}\lambda_{1}(\Sigma_{w})}{1-\delta_{x}^{\frac{1}{tn_{x}}}}}$
with probability at least $\delta_{x}$. Given the probability $\delta_{x}$, the upper bound of the sample $\hat{C}_{k}$ is given by:
(84) $\begin{split}&~{}~{}~{}~{}\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~{}\top}(Q+\hat{K}^{\top}_{k}R\hat{K}_{k})x_{t}^{(k)}\\ &\leq\frac{1}{l}\sum_{t=0}^{l-1}\lVert Q_{\hat{K}_{k}}\rVert\lVert x_{t}^{(k)}\rVert^{2}\leq(\lVert Q_{\hat{K}_{k}}\rVert\max_{t\in[0,l-1]}\lVert x_{t}^{(k)}\rVert^{2}\leq\frac{C(\hat{K}_{k})}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)L_{w}\right)^{2}\\ &\leq\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\sqrt{\frac{n_{x}\lambda_{1}(\Sigma_{w})}{1-\delta_{x}^{\frac{1}{ln_{x}}}}}\right)^{2}=:\bar{C}\end{split}$
Each sample $Z_{k}:=\frac{n_{x}n_{u}}{r^{2}}\hat{C}_{k}U_{k}$ has bounded norm as:
(85) $\lVert Z_{k}\rVert\leq\frac{n_{x}n_{u}}{r}\bar{C}=:\alpha_{6}(C(K),\epsilon_{x}).$
@@
(86) $\lVert Z\rVert\leq\epsilon_{l}+\lVert\bar{\nabla}C(K)\rVert\leq\epsilon_{l}+\epsilon_{nr}+\lVert\nabla{C}(K)\rVert\leq\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))$
@@
(87) $\lVert Z_{k}-Z\rVert\leq\lVert Z\rVert+\lVert Z_{k}\rVert\leq\underbrace{\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))+\alpha_{6}(C(K),\delta_{x})}_{=:\alpha_{7}(C(K),\delta_{x})}$
@@
(88) $\begin{split}\lVert\mathbb{E}(Z_{k}Z_{k}^{\top})-ZZ^{\top}\rVert&\leq\lVert\mathbb{E}(Z_{k}^{\top}Z_{k})\rVert_{F}+\lVert Z^{\top}Z\rVert_{F}\leq\max_{Z_{k}}(\lVert Z_{k}\rVert_{F})^{2}+\lVert Z\rVert^{2}_{F}\\ &\leq\underbrace{\max\{n_{x},n_{u}\}^{2}\alpha^{2}_{6}(C(K),\epsilon_{x})+\left(\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))\right)^{2}}_{=:\alpha_{8}(C(K),\delta_{x})}\end{split}$
Given an accuracy $\epsilon_{d}$ and $\delta_{d}$, we can use the matrix concentration inequality [11, Lemma B2.5],
(89) $N=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{d}^{2}}\left(\alpha_{7}^{2}(C(K),\delta_{x})+\frac{\alpha_{8}(C(K),\delta_{x})\epsilon_{d}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{d}}\right]$
@@
(90) [^$\mathbb{P}\left[\lVert\hat{\nabla}C(K)-\nabla^{\prime}C(K)\rVert_{F}\leq\epsilon_{d}\right]\geq 1-\delta_{d}.$
Now, we can summarize all the terms. Given an arbitrary tolerance $\epsilon$, which can be decomposed by $\epsilon_{d},\epsilon_{l},\epsilon_{n},\epsilon_{r}$, with $\epsilon=\epsilon_{d}+\epsilon_{l}+\epsilon_{n}+\epsilon_{r}$, and an arbitrary probability $\delta$, which can be decomposed by $\delta_{x},\delta_{d},\delta_{n}$ with $\delta=(1-\delta_{d})(1-\delta_{n})\delta_{x}$, suppose the exploration radius is chosen as $r\leq r_{\max}$ with
(91) $r_{\max}=\min\{h(C(K)),\lVert K\rVert,\frac{\epsilon_{r}}{h_{\nabla}(C(K))}\},$
@@
(92) $l_{\min}=\frac{2(C(K)+rh_{C}C(K))^{2}}{\epsilon_{l}\lambda_{1}(\Sigma_{w})}\left(\frac{\lVert\Sigma_{0}\rVert\lambda_{1}(\Sigma_{w})+C(K)+rh_{C}C(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{1}{\lambda_{1}(Q)}\right),$
and the number of rollouts is chosen as: $n\geq N$, where $N:=\max\{N_{1},N_{2}\}$ with
(93a) $N_{1}$ $:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{n}^{2}}\left(\alpha_{4}^{2}(C(K),\epsilon_{r})+\frac{\alpha_{5}(C(K),\epsilon_{r})\epsilon_{n}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{n}}\right],$
(93b) $N_{2}$ $:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{d}^{2}}\left(\alpha_{7}^{2}(C(K),\delta_{x})+\frac{\alpha_{8}(C(K),\delta_{x})\epsilon_{d}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{d}}\right],$
@@
{^$\mathbb{P}\left\{\lVert\hat{\nabla}C(K)-{\nabla}C(K)\rVert\leq\epsilon\right\}\geq\delta.$
@@ C.6 Proof of Theorem 31
(94) $\hat{\Sigma}_{K}-\Sigma_{K}=(\hat{\Sigma}_{K}-\tilde{\Sigma}^{(l)}_{K})+(\tilde{\Sigma}^{(l)}_{K}-\tilde{\Sigma}_{K})+(\tilde{\Sigma}_{K}-\Sigma_{K}),$
@@
For the last term, we aim to bound $\lVert\tilde{\Sigma}_{K}-\Sigma_{K}\rVert$. From Lemma 2.2, if we choose $r\leq r^{\prime}_{\max}$ as $r^{\prime}_{\max}:=\frac{\epsilon^{\prime}_{r}}{b_{\nabla}(C(K))}$, where $b_{\nabla}$ is defined in (56), then $\lVert\tilde{\Sigma}_{K}-\Sigma_{K}\rVert\leq{\epsilon^{\prime}_{r}}$.
For the second term, from Lemma C.1, if we choose $l\geq l^{\prime}_{\min}$ with
@@
For any desired probability $\delta^{\prime}_{x}$, $\lVert x_{t}^{(k)}\rVert$ is upper-bounded by $\bar{L}^{\prime}$ which is defined as $\bar{L}^{\prime}:=L_{0}+(l-1)\sqrt{\frac{n_{x}\lambda_{1}(\Sigma_{w})}{1-(\delta_{x}^{\prime})^{\frac{1}{ln_{x}}}}}$(see Lemma 76). Because of the definition of $\tilde{\Sigma}^{(l)}_{K}$, we have:
@@
(95) $\lVert\tilde{Z}_{k}-\tilde{Z}\rVert\leq\lVert\tilde{Z}_{k}\rVert+\lVert\tilde{Z}\rVert\leq\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(Q)}+(\bar{L}^{\prime}(\delta^{\prime}_{x}))^{2}=:\alpha_{9}(C(K),\delta^{\prime}_{x})$
@@
(96) $\begin{split}\lVert\mathbb{E}(\tilde{Z}_{k}\tilde{Z}_{k}^{\top})-\tilde{Z}\tilde{Z}^{\top}\rVert&\leq\lVert\mathbb{E}(\tilde{Z}_{k}^{\top}\tilde{Z}_{k})\rVert_{F}+\lVert\tilde{Z}^{\top}\tilde{Z}\rVert_{F}\\ &\leq\max_{\tilde{Z}_{k}}(\lVert\tilde{Z}_{k}\rVert_{F})^{2}+\lVert\tilde{Z}\rVert^{2}_{F}\\ &\leq\underbrace{n_{x}^{2}\left[(\bar{L}^{\prime}(\delta^{\prime}_{x}))^{2}+\left(\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(Q)}\right)^{2}\right]}_{=:\alpha_{10}(C(K),\delta^{\prime}_{x})}\end{split}$
With Matrix Bernstein Inequality [11, Lemma B2.5], given a small tolerance ${\epsilon^{\prime}_{n}}$ and small probability $\delta^{\prime}_{n}$, if the number of rollouts $n\geq n_{\min}^{\prime}$ with
@@
We sum all terms, we can derive that given an arbitrary tolerance $\epsilon^{\prime}$, which can be decomposed by $\epsilon^{\prime}_{n},\epsilon^{\prime}_{l},\epsilon^{\prime}_{r}$, with $\epsilon=\epsilon^{\prime}_{n}+\epsilon^{\prime}_{l}+\epsilon^{\prime}_{r}$, and an arbitrary probability $\delta^{\prime}$, which can be decomposed by $\delta^{\prime}_{x},\delta^{\prime}_{n}$ with $\delta=(1-\delta_{n}^{\prime})\delta_{x}$, suppose the exploration radius is chosen as $r\leq r^{\prime}_{\max}$ with
(97) $r^{\prime}_{\max}:=\min\left\{h(C(K)),\lVert K\rVert,\frac{\epsilon^{\prime}_{r}}{b_{\nabla}(C(K))}\right\}$
@@
(98) $l^{\prime}_{\min}:=\frac{2C(K)}{\epsilon^{\prime}_{l}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right)$
@@
(99) $n_{\min}^{\prime}:=\frac{2n_{x}}{(\epsilon^{\prime}_{n})^{2}}\left(\alpha_{10}(C(K),\delta^{\prime}_{x})+\frac{\alpha_{9}(C(K),\delta^{\prime}_{x})\epsilon^{\prime}_{n}}{3\sqrt{n_{x}}}\right)\log\left[\frac{2n_{x}}{\delta^{\prime}_{n}}\right]$
@@
$\mathbb{P}\left\{\lVert\hat{\Sigma}_{K}-\Sigma_{K}\rVert\leq\epsilon^{\prime}\right\}\geq\delta^{\prime}.$

### v2

@@ C.5 Proof of Theorem 4
$\hat{\nabla}C(K)=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}\hat{C}_{\bar{K}_{k}}U_{k}$
and the true gradient $\nabla C(K)$.
Proof 11
We bridge this two terms with $\nabla^{\prime}C(K)$ and $\bar{\nabla}C(K))$, where $\bar{\nabla}C(K))$ is defined in Lemma 8 and $\nabla^{\prime}C(K)$ is defined as:
$\nabla^{\prime}C(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}C^{(l)}(K+U_{k})U_{k},$
where $C^{(l)}(K)$ is defined in Lemma 6. The relation between the estimates and the true gradient is given as:
$\hat{\nabla}C(K)-\nabla C(K)=(\hat{\nabla}C(K)-\nabla^{\prime}C(K))+(\nabla^{\prime}C(K)-\bar{\nabla}C(K))+(\bar{\nabla}C(K)-\nabla C(K))$ (78)
The third term is analyzed in Lemma 8. We choose $\epsilon_{nr}:=\epsilon_{n}+\epsilon_{r}$ and $\delta_{n}$, then we can find the exploration radius $r_{\max}$ and $N_{1}$ as:
$r_{\max}$ $(C(K),\epsilon_{r})=\min\{h(C(K)),\lVert K\rVert,\frac{\epsilon_{r}}{h_{\nabla}(C(K))}\},$ (79a)
$N_{1}$ $=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{n}^{2}}\left(\alpha_{4}^{2}(C(K),\epsilon_{r})+\frac{\alpha_{5}(C(K),\epsilon_{r})\epsilon_{n}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{n}}\right]$ (79b)
@@
$\mathbb{P}\left[\lVert\bar{\nabla}C(K)-\nabla C(K)\rVert_{F}\leq\epsilon_{nr}\right]\geq 1-\delta_{n}.$ (80)
For the second term, as stated in Lemma 6, for any desired accuracy $\epsilon_{l}$, we can choose $l$ such that:
$l\geq\frac{2n_{x}n_{u}(C(K)+rh_{C}C(K))^{2}}{\epsilon_{l}r\lambda_{1}(\Sigma_{w})}\left(\frac{\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{(C(K)+rh_{C}C(K))}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{1}{\lambda_{1}(Q)}\right),$ (81)
@@
$|C^{(l)}(\bar{K}_{k})-C(\bar{K}_{k})|\leq\frac{r\epsilon_{l}}{n_{x}n_{u}},\quad\forall k\in[1,n].$ (82)
@@
$\lVert\nabla^{\prime}C(K)-\bar{\nabla}C(K)\rVert\leq\left\lVert\frac{1}{n}\sum_{k=1}^{n}\frac{n_{x}n_{u}}{r^{2}}\left[C^{(l)}(\bar{K}_{k})-C(\bar{K}_{k})\right]U_{k}\right\rVert=\epsilon_{l}$ (83)
@@
$\lVert x_{t}\rVert\leq\bar{L}=L_{0}+t\bar{w}=L_{0}+t\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/t}}$ (84)
with probability at least $1-\delta_{x}$. Given the probability $\delta_{x}$, the upper bound of the sample $\hat{C}_{\bar{K}_{k}}$ is given by:
$\begin{split}&~~~~\frac{1}{l}\sum_{t=0}^{l-1}x_{t}^{(k)~\top}(Q+\bar{K}^{\top}_{k}R\bar{K}_{k})x_{t}^{(k)}\\ &\leq\frac{1}{l}\sum_{t=0}^{l-1}\lVert Q_{\bar{K}_{k}}\rVert\lVert x_{t}^{(k)}\rVert^{2}\leq(\lVert Q_{\bar{K}_{k}}\rVert\max_{t\in[0,l-1]}\lVert x_{t}^{(k)}\rVert^{2})\leq\frac{C(\bar{K}_{k})}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\bar{L}\right)^{2}\\ &\leq\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/l}}\right)^{2}=:\bar{C}\end{split}$ (85)
Each sample $Z_{k}:=\frac{n_{x}n_{u}}{r^{2}}\hat{C}_{\bar{K}_{k}}U_{k}$ has bounded norm as:
$\lVert Z_{k}\rVert\leq\frac{n_{x}n_{u}}{r}\bar{C}=:\alpha_{6}(C(K),\delta_{x}).$ (86)
@@
$\lVert Z\rVert\leq\epsilon_{l}+\lVert\bar{\nabla}C(K)\rVert\leq\epsilon_{l}+\epsilon_{nr}+\lVert\nabla{C}(K)\rVert\leq\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))$ (87)
@@
$\lVert Z_{k}-Z\rVert\leq\lVert Z\rVert+\lVert Z_{k}\rVert\leq\underbrace{\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))+\alpha_{6}(C(K),\delta_{x})}_{=:\alpha_{7}(C(K),\delta_{x})}$ (88)
@@
$\begin{split}\lVert\mathbb{E}(Z_{k}Z_{k}^{\top})-ZZ^{\top}\rVert&\leq\lVert\mathbb{E}(Z_{k}^{\top}Z_{k})\rVert_{F}+\lVert Z^{\top}Z\rVert_{F}\leq\max_{Z_{k}}(\lVert Z_{k}\rVert_{F})^{2}+\lVert Z\rVert^{2}_{F}\\ &\leq\underbrace{\max\{n_{x},n_{u}\}^{2}\alpha^{2}_{6}(C(K),\epsilon_{x})+\left(\epsilon_{l}+\epsilon_{nr}+b_{\nabla}(C(K))\right)^{2}}_{=:\alpha_{8}(C(K),\delta_{x})}\end{split}$ (89)
Given an accuracy $\epsilon_{d}$ and $\delta_{d}$, we can use the matrix concentration inequality [10, Lemma B2.5],
$N_{2}=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{d}^{2}}\left(\alpha_{7}^{2}(C(K),\delta_{x})+\frac{\alpha_{8}(C(K),\delta_{x})\epsilon_{d}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{d}}\right]$ (90)
@@
$\mathbb{P}\left[\lVert\hat{\nabla}C(K)-\nabla^{\prime}C(K)\rVert_{F}\leq\epsilon_{d}\right]\geq 1-\delta_{d}.$ (91)
Now, we can summarize all the terms. Given an arbitrary tolerance $\epsilon$, which can be expressed by $\epsilon_{d},\epsilon_{l},\epsilon_{n},\epsilon_{r}$, with $\epsilon=\epsilon_{d}+\epsilon_{l}+\epsilon_{n}+\epsilon_{r}$, and an arbitrary probability $\delta$, which can be expressed by $\delta_{x},\delta_{d},\delta_{n}$ with $1-\delta=(1-\delta_{d})(1-\delta_{n})(1-\delta_{x})$, suppose the exploration radius is chosen as $r\leq r_{\max}$ with
$r_{\max}=\min\{h(C(K)),\lVert K\rVert,\frac{\epsilon_{r}}{h_{\nabla}(C(K))}\},$ (92)
@@
$l_{\min}=\frac{2n_{x}n_{u}(C(K)+rh_{C}C(K))^{2}}{\epsilon_{l}r\lambda_{1}(\Sigma_{w})}\left(\frac{\lVert\Sigma_{0}\rVert\lambda_{1}(\Sigma_{w})+C(K)+rh_{C}C(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{1}{\lambda_{1}(Q)}\right),$ (93)
and the number of rollouts is chosen as: $n\geq\max\{N_{1},N_{2}\}$ with
$N_{1}$ $:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{n}^{2}}\left(\alpha_{4}^{2}(C(K),\epsilon_{r})+\frac{\alpha_{5}(C(K),\epsilon_{r})\epsilon_{n}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{n}}\right],$ (94a)
$N_{2}$ $:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{d}^{2}}\left(\alpha_{7}^{2}(C(K),\delta_{x})+\frac{\alpha_{8}(C(K),\delta_{x})\epsilon_{d}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{d}}\right],$ (94b)
@@
$\mathbb{P}\left\{\lVert\hat{\nabla}C(K)-{\nabla}C(K)\rVert\leq\epsilon\right\}\geq 1-\delta.$
@@ C.6 Proof of Theorem 5
$\hat{\Sigma}_{K}-\Sigma_{K}=(\hat{\Sigma}_{K}-\tilde{\Sigma}^{(l)}_{K})+(\tilde{\Sigma}^{(l)}_{K}-\tilde{\Sigma}_{K})+(\tilde{\Sigma}_{K}-\Sigma_{K}),$ (95)
@@
For the last term, we aim to bound $\lVert\tilde{\Sigma}_{K}-\Sigma_{K}\rVert$. From Lemma 3, if we choose $r\leq r^{\prime}_{\max}$ as $r^{\prime}_{\max}:=\frac{\epsilon^{\prime}_{r}}{b_{\nabla}(C(K))}$, where $b_{\nabla}$ is defined in (56), then $\lVert\tilde{\Sigma}_{K}-\Sigma_{K}\rVert\leq{\epsilon^{\prime}_{r}}$.
For the second term, from Lemma 6, if we choose $l\geq l^{\prime}_{\min}$ with
@@
For any desired probability $\delta^{\prime}_{x}$, $\lVert x_{t}^{(k)}\rVert$ is upper-bounded by $\bar{L}^{\prime}$ which is defined as $\bar{L}^{\prime}(\delta^{\prime}_{x}):=\left(L_{0}+t\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta^{\prime}_{x})^{1/t}}\right)$(see Lemma 9). Because of the definition of $\tilde{\Sigma}^{(l)}_{K}$, we have:
@@
$\lVert\tilde{Z}_{k}-\tilde{Z}\rVert\leq\lVert\tilde{Z}_{k}\rVert+\lVert\tilde{Z}\rVert\leq\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(Q)}+(\bar{L}^{\prime}(\delta^{\prime}_{x}))^{2}=:\alpha_{9}(C(K),\delta^{\prime}_{x})$ (96)
@@
$\begin{split}\lVert\mathbb{E}(\tilde{Z}_{k}\tilde{Z}_{k}^{\top})-\tilde{Z}\tilde{Z}^{\top}\rVert&\leq\lVert\mathbb{E}(\tilde{Z}_{k}^{\top}\tilde{Z}_{k})\rVert_{F}+\lVert\tilde{Z}^{\top}\tilde{Z}\rVert_{F}\\ &\leq\max_{\tilde{Z}_{k}}(\lVert\tilde{Z}_{k}\rVert_{F})^{2}+\lVert\tilde{Z}\rVert^{2}_{F}\\ &\leq\underbrace{n_{x}^{2}\left[(\bar{L}^{\prime}(\delta^{\prime}_{x}))^{2}+\left(\frac{C(K)+rh_{C}(C(K))}{\lambda_{1}(Q)}\right)^{2}\right]}_{=:\alpha_{10}(C(K),\delta^{\prime}_{x})}\end{split}$ (97)
With Matrix Bernstein Inequality [10, Lemma B2.5], given a small tolerance $\epsilon^{\prime}_{n}$} and small probability $\delta^{\prime}_{n}$, if the number of rollouts $n\geq n_{\min}^{\prime}$ with
@@
We sum all terms, we can derive that given an arbitrary tolerance $\epsilon^{\prime}$, which can be expressed by $\epsilon^{\prime}_{n},\epsilon^{\prime}_{l},\epsilon^{\prime}_{r}$, with $\epsilon=\epsilon^{\prime}_{n}+\epsilon^{\prime}_{l}+\epsilon^{\prime}_{r}$, and an arbitrary probability $\delta^{\prime}$, which can be expressed by $\delta^{\prime}_{x},\delta^{\prime}_{n}$ with $1-\delta^{\prime}=(1-\delta_{n}^{\prime})(1-\delta_{x}^{\prime})$, suppose the exploration radius is chosen as $r\leq r^{\prime}_{\max}$ with
$r^{\prime}_{\max}:=\min\left\{h(C(K)),\lVert K\rVert,\frac{\epsilon^{\prime}_{r}}{b_{\nabla}(C(K))}\right\}$ (98)
@@
$l^{\prime}_{\min}:=\frac{2C(K)}{\epsilon^{\prime}_{l}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right)$ (99)
@@
$n_{\min}^{\prime}:=\frac{2n_{x}}{(\epsilon^{\prime}_{n})^{2}}\left(\alpha_{10}(C(K),\delta^{\prime}_{x})+\frac{\alpha_{9}(C(K),\delta^{\prime}_{x})\epsilon^{\prime}_{n}}{3\sqrt{n_{x}}}\right)\log\left[\frac{2n_{x}}{\delta^{\prime}_{n}}\right]$ (100)
@@
$\mathbb{P}\left\{\lVert\hat{\Sigma}_{K}-\Sigma_{K}\rVert\leq\epsilon^{\prime}\right\}\geq 1-\delta^{\prime}.$